OpenStudy (anonymous):

a rectangular park 80 meters long and 60 meters wide is sorrounded by a driveway of uniform width. how wide is the drive way if its area is 3/25 of the area of the park? answer pls

OpenStudy (anonymous):

pls help i need it for my project :(

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{green}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{green}{Welcome~to~\color{green}{Open}}\color{green}{Study!!!!}\) Hello poleng. We are not going to help if we don't want to. It's our choice. But I'm that nice so I'll help you out a little:) Let x = width of driveway. 80 + 2x = length including driveway 60 + 2x = width including driveway (80 + 2x)(60 + 2x) = 80(60) + (3/25)(80)(60) 4800 + 280x + 4x^2 = 4800 + 576 4x^2 + 280x - 576 = 0 x^2 + 70x - 144 = 0 (x + 72)(x - 2) = 0 (x + 72) = 0 x = -72 Discard, negative (x - 2) = 0 x = 2 meter, width of driveway

OpenStudy (anonymous):

thank you for your kindness YanaSidlinkskiy :)

OpenStudy (yanasidlinskiy):

You're most Welcome!!!!:) Enjoy OpenStudy!!:)

OpenStudy (anonymous):

:)

OpenStudy (yanasidlinskiy):

Also if your question has been answered, please close the post and open up new questions(if you have any)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

ok ill post another question :)